A Binomial Experiment has 5 trials. Each trial has a probability of success of .7. Compute the probability of having exactly 2 successes. Your final answer should be correct to 3 places after the decimal point.
Answer #1

solution

Given that ,

p = 0.7

1 - p = 1 - 0.7 = 0.3

n =5

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 2) = (5 C 2) * 0.72 * (0.3)3

probability =0.132